# ReaGAN: Node-as-Agent-Reasoning Graph Agentic Network

### v1

@@ 1 Introduction
To address the two key challenges, we propose a new perspective on graph learning by rethinking the fundamental computational unit: we treat each node as an autonomous agent. Unlike GNNs that passively aggregate messages through static rules, each node acts as an agent who actively determines its action at each layer based on internal states and contextual signals. Inspired by general agent-based systems (Russell and Norvig, 2010; Wang et al., 2023; Xu et al., 2025), our agentic framework endows each node with four core components: Planning, which decides the next operations based on its current objective and context; Actions, which executes local-global aggregation strategies to interact with neighbors (Yao et al., 2022; Shinn et al., 2023); Memory, which stores the node’s cumulative textual features along with historical interaction traces; and Tool Use, which calls external functions like retrieval-augmented generation. In this work, we unify these components into the Retrieval-augmented Graph Agentic Network (ReaGAN), which enables each node to make individualized, adaptive, and semantically informed decisions. This self-decision design directly addresses Challenge 1 by granting each node autonomy in its message passing, while the local-global aggregation strategy addresses Challenge 2 as the following paragraph shows.
Beyond Local Aggregation which enables nodes to exchange messages with directly connected neighbors, we integrate retrieval-augmented generation (RAG) (Lewis et al., 2020) as an external tool that empowers nodes to access semantically relevant but structurally distant information, where the entire graph is viewed as a searchable database. Each node then integrates both local structural signals and globally retrieved semantic content into its individual memory, forming a richer contextual representation for subsequent planning and interaction. Equipped with these capabilities, each node engages in an agentic workflow. Based on its current state and memory, the node prompts a frozen LLM to generate a context-aware plan, selects and executes the appropriate actions, such as local or global aggregation, optionally invokes external tools like RAG, and subsequently updates its internal memory with the acquired information.
@@ 2.2 Action: Node-Level Decision Space
[…]
@@
[…]
@@ 2.4 Tools: Global Semantic Augmentation
[…]
@@
[…]
@@ 6 Limitations and Future Work
In future work, we envision ReaGAN as a foundational layer in multi-agent systems (Li et al., 2025), where each node-agent can participate in decentralized decision-making(Huang et al., 2025) or inter-agent communication. For instance, its ability to reason locally and retrieve information globally makes it a strong fit for modular systems (Gershon et al., 2025) or routing-based architectures (Mei et al., 2025a), where nodes may act as autonomous modules coordinating across a shared substrate. Additionally, to scale ReaGAN under constrained resources, we propose integrating it with agent orchestration frameworks such as AIOS (Mei et al., 2025b)—enabling parallelized, resource-aware scheduling of reasoning agents across large graphs. Rather than being limited to graph classification, we see ReaGAN as a general blueprint for scalable agentic inference—well suited for future systems that demand fine-grained, context-aware, and communication-capable reasoning units.

### v2

@@ 1 Introduction
To address the two key challenges, we propose a new perspective on graph learning by rethinking the fundamental computational unit: we treat each node as an autonomous agent. Unlike GNNs that passively aggregate messages through static rules, each node acts as an agent who actively determines its action at each layer based on internal states and contextual signals. Inspired by general agent-based systems (Russell and Norvig, 2010; Wang et al., 2023; Xu et al., 2025; Sun et al., ; Gao and Zhang, 2024), our agentic framework endows each node with four core components: Planning, which decides the next operations based on its current objective and context; Actions, which executes local-global aggregation strategies to interact with neighbors (Yao et al., 2022; Shinn et al., 2023); Memory, which stores the node’s cumulative textual features along with historical interaction traces; and Tool Use, which calls external functions like retrieval-augmented generation. In this work, we unify these components into the Retrieval-augmented Graph Agentic Network (ReaGAN), which enables each node to make individualized, adaptive, and semantically informed decisions. This self-decision design directly addresses Challenge 1 by granting each node autonomy in its message passing, while the local-global aggregation strategy addresses Challenge 2 as the following paragraph shows.
Beyond Local Aggregation which enables nodes to exchange messages with directly connected neighbors, we integrate retrieval-augmented generation (RAG) (Lewis et al., 2020; Su et al., 2025) as an external tool that empowers nodes to access semantically relevant but structurally distant information, where the entire graph is viewed as a searchable database. Each node then integrates both local structural signals and globally retrieved semantic content into its individual memory, forming a richer contextual representation for subsequent planning and interaction. Equipped with these capabilities, each node engages in an agentic workflow. Based on its current state and memory, the node prompts a frozen LLM to generate a context-aware plan, selects and executes the appropriate actions, such as local or global aggregation, optionally invokes external tools like RAG, and subsequently updates its internal memory with the acquired information.
@@ 2.2 Action: Node-Level Decision Space
    where $\texttt{TextAgg}(\cdot)$ denotes a natural language-level aggregation function such as concatenation or summarization. Here, $\tilde{t}_{v}^{(l-1)}$ is the node’s aggregated feature from the previous layer (initialized as $t_{v}$). The result $\tilde{t}_{v}^{(l)}$ is stored in memory as an updated representation of the node’s semantic environment.
@@
The Global Aggregation action enables a node to augment its semantic context by retrieving structurally distant but semantically similar nodes from across the graph. This is achieved by invoking a retrieval tool (e.g., RAG), which queries a structure-free database constructed from all node-level textual representations. Unlike the original graph, this database preserves no edges; each entry consists of a node’s text feature $t_{u}$ and its label $y_{u}$ (if available). Retrieval is based purely on textual semantic similarity, typically computed via dense embedding similarity (e.g., cosine distance). Given a query node’s current textual state $t_{v}^{(l-1)}$, the system retrieves the Top-$K$ most relevant global semantic neighbors:
@@ 2.4 Tools: Global Semantic Augmentation
To support global semantic reasoning, ReaGAN equips each node with a single external tool: Retrieval-Augmented Generation (RAG). This tool enables a node to retrieve semantically relevant information from across the entire graph—beyond its structural neighborhood. A structure-free database is constructed, consisting of all nodes’ text features and, when available, their labels. Each node contributes a single entry, and the resulting corpus $\mathcal{D}$ is indexed by textual similarity. Given a node’s current representation $t_{v}^{(l)}$, the RAG tool performs a top-$K$ similarity search over the database:
@@
where $\texttt{sim}(\cdot,\cdot)$ denotes an embedding-based similarity function (e.g., cosine distance).
@@ 6 Limitations and Future Work
In future work, we envision ReaGAN as a foundational layer in multi-agent systems (Li et al., 2025), where each node-agent can participate in decentralized decision-making(Huang et al., 2025) or inter-agent communication. For instance, its ability to reason locally and retrieve information globally makes it a strong fit for modular systems (Gershon et al., 2025; Jin et al., 2025) or routing-based architectures (Mei et al., 2025a), where nodes may act as autonomous modules coordinating across a shared substrate. Additionally, to scale ReaGAN under constrained resources, we propose integrating it with agent orchestration frameworks such as AIOS (Mei et al., 2025b)—enabling parallelized, resource-aware scheduling of reasoning agents across large graphs. Rather than being limited to graph classification, we see ReaGAN as a general blueprint for scalable agentic inference—well suited for future systems that demand fine-grained, context-aware, and communication-capable reasoning units.